OpenStudy (anonymous):

A tennis ball is thrown with an initial upward velocity of 64 feet per second. How long will it take for the ball to hit the ground? Use the formula in which the height above ground after t seconds is given by the quadratic function h = vt – 16t2, where h represents the height of the object in feet, and v represents the object's initial velocity in feet per second.

OpenStudy (anonymous):

@Loujoelou ? :p

OpenStudy (anonymous):

We aren't solving for time right. just the h?

OpenStudy (anonymous):

I guess h lol

OpenStudy (anonymous):

A:3 seconds C:-3 seconds B:-4 seconds D:4 seconds

OpenStudy (anonymous):

K. -16t^2+64+0 t=-b/2a So we will solve for t 64/(-16)(2) 64/-32 t=-2 Now we will put this back into the equation and solve for h -16(-2)^2+64(-2)+0 and we get h=-192. I'm guessing maybe we have to do -192/64=-3? Not sure about solving for the seconds sry.